# Topological signatures of collective dynamics and turbulent-like energy cascades in active granular matter

### v1

@@ Introduction
To explore this physical mechanism further, we calculate $D_{\text{min}}^{2}$, which quantifies the strength of nonaffine particle displacements between two consecutive configurations (see Methods A for details) [74]. A large value for this quantity represents soft areas in the sense that structural relaxation is more intense than elsewhere. In Fig.3(a), we calculate the intensity of $D_{\text{min}}^{2}$ projected on each particle for $\tau=500s$ using a gray color map, with darker regions representing larger values of $D_{\text{min}}^{2}$. By visually comparing this color map with that presented in the third panel of Fig.2(a), it is clear that regions with large $D_{\text{min}}^{2}$ correspond to smaller values of $P_{c}$. This implies that areas with strong structural relaxation are those where motion is still random and poorly collective.
In Fig. 3(a), we represent the topological defects (TDs) with charges $+1$ and $-1$ using blue and red symbols, respectively. These TDs appear in pairs within regions of higher $D_{\text{min}}^{2}$ values. To quantify this observation, we define ‘soft spots’ (SS) as the top 5% of particles with the highest $D_{\text{min}}^{2}$ values and analyze their positional correlation with both positive and negative defects. The results are shown in Fig. 3(b) and (c). Each curve exhibits a peak at small $r$, with the peak height increasing slightly over time. This suggests a strong short-range correlation between soft spots and TDs. The correlation is slightly stronger for negative defects, though the difference is not substantial. This could be explained by the fact that TDs typically appear in pairs, which may balance out the correlation differences. Overall, the observations in Fig. 3 suggest that topological defects play a key role in the system’s structural relaxation and mediate the emergence of collective motion during out-of-equilibrium dynamics.
@@
Figure 5: Turbulent-like energy cascades as revealed by energy power spectra. Longitudinal (a) and transverse (b) kinetic energy spectra at different observation times $\tau$ as a function of the wave number $k$. (c) Power-law exponents $b$ for longitudinal ($b_{L}$) and transverse spectra ($b_{T}$).
@@
We investigate the non-equilibrium dynamics of the system across a full range of observation times, spanning six decades from $\tau=10^{-2}\,s$ to $\tau=10^{4}\,s$, and reveal a three-stage evolution that is evident in both topological characteristics, particle level dynamics, and collective motion. Three snapshots of the system dynamics in each of these stages have already been presented in Fig. 2 (see vertical red arrows in Fig. 4).
@@
Figure 4(b) displays the mean square displacement (MSD), where the unit length is given in terms of the small particle diameter ($d_{s}$). This quantity characterizes the system’s dynamics at the particle level. The MSD evolution follows three distinct stages. For $t<t_{1}$, the MSD exhibits sub-diffusive behavior, $\text{MSD}\propto t^{0.32}$, indicating that particle motion is constrained within a ‘cage’ on short timescales. Between $t_{1}$ and $t_{2}$, as the total number of TDs, $N_{\text{defects}}$, decreases more rapidly, particle dynamics accelerate while maintaining sub-diffusive characteristics with a different exponent, $\text{MSD}\sim t^{0.55}$. This regime coincides with the increasing mean cluster size of collective motion, as observed in the second snapshot of Fig. 2(a). Beyond $t_{2}$, particle dynamics slow as the system reaches a non-equilibrium steady state, where the MSD approaches a plateau. This trend aligns with the stabilization of the total number of defects, confirming that defects mediate local relaxation. Ultimately, the system attains a collective steady state, emerging from the interplay between activity (energy injected through vibrations) and inelastic collisions.
@@
In conclusion, the number of defects, MSD, and average curl exhibit three distinct dynamical stages, establishing a link between topology, particle-level dynamics, and collective motion. These quantities are interrelated: at short time scales ($t<t_{1}$), particle motion is “trapped” due to frequent collisions, resulting in slow sub-diffusion in the MSD. Rapid fluctuations in the displacement field, driven by collisions, lead to efficient defect annihilation, causing a continuous decrease in defect numbers. Between $t_{1}$ and $t_{2}$, the system’s dynamics evolve more rapidly, driven by the formation of larger vortices in the displacement field. During this period, the defect number decreases at an accelerated rate. After $t_{2}$, the system reaches a steady state, where the number of defects, MSD, and average curl stabilize over time. The average curl attains its maximum value, signifying the emergence of stable large vortices. This accounts for the total topological number approaching a plateau and the onset of large-scale collective motion. Defects regulate local structural rearrangements, ultimately culminating in the emergence of large vortices over long timescales.
@@
Nevertheless, a clear trend emerges in the intermediate and low range of $k$. In particular, at short times, both $E_{L}$ and $E_{T}$ grow with $k$, flattening in an intermediate $\tau$ range and then changing their slope at late times. To better characterize this behavior, we fit the various curves with a power-law ansatz $E_{L,T}\propto k^{-b_{L,T}}$. The extracted power-law exponent is presented in Fig. 5(c) as a function of the observation time $\tau$. Below $t_{1}$, we observe that $b_{L,T}$ are approximately identical and both negative, signaling a spectrum that grows towards smaller scales. Around $t_{1}$, the spectrum starts to flatten, and the power-law exponents become slightly positive, presenting differences between the longitudinal and transverse sectors. Finally, above $t_{2}$, the power-law exponents reach a plateau value in approximately $b_{L}\approx 0.3$ and $b_{T}\approx 0.4$.
This kinematic behavior aligns well with the three-stage dynamics presented in Fig. 4 and can be rationalized as follows. At short time scales, the energy injected by the active driving is gradually transferred to the individual particles, inducing a direct energy cascade towards microscopic scales. Around $t_{1}\approx 3\,s$, energy is evenly distributed across all scales, corresponding to the highest degree of dynamical heterogeneities. When the observation time exceeds $3\,s$, energy starts transferring from small to larger scales, exhibiting a turbulent-like inverse cascade. This process becomes steady around $t_{2}\approx 60\,s$ when the energy transfer becomes roughly independent of time.
@@
We assume that the mean squared vorticity, $\overline{\Omega^{2}}$, follows $\overline{\Omega^{2}}\propto(\Omega_{0}/\tau^{\alpha})^{2}$. By definition, this implies $\overline{\Omega^{2}}\propto\tau^{-2\alpha}\textbf{u}_{l}^{2}$ from dimensional analysis. Estimating $\textbf{u}_{l}^{2}$ from the MSD results in Fig. 4(b), we obtain $\textbf{u}_{l}^{2}\propto\tau^{\beta}$, with $\beta=0.3$ for $\tau<t_{1}$ and $\beta=0.55$ for $t_{1}<\tau<t_{2}$. This leads to $\overline{\Omega^{2}}\propto\tau^{\beta-2\alpha}$. Notably, the exponent $\beta-2\alpha\approx 0$ for both $\tau<t_{1}$ and $t_{1}<\tau<t_{2}$. These findings suggest that the mean squared vorticity is nearly conserved in the turbulent-like inverse energy cascade of our active granular system, adhering to a principle akin to that of 2D turbulence.
@@
Our initial analysis focused on the highest packing fraction ($\phi=0.822$), where strong inelastic collisions balance the effects of activity, resulting in a late-time non-equilibrium steady state. Here, we extend the analysis to the lowest packing fraction ($\phi=0.309$), where the dynamics remain decorrelated and gas-like due to the weak inelastic interactions [38].
Figure 6(a) shows the evolution of the total number of topological defects, $N_{\text{defects}}$, for four packing fractions ($\phi=0.309$, $0.618$, $0.773$, and $0.811$). At a high packing fraction ($\phi=0.811$), the behavior of $N_{\text{defects}}$ resembles that observed at the maximal packing fraction shown in Fig. 4(a).
@@ Conclusions
Using active granular disks as a model for nonpolar dry active Brownian particles, we conducted a comprehensive experimental study to investigate the topological features that give rise to steady-state large-scale collective dynamics and turbulent-like inverse energy cascades. The dynamical behavior of the system evolves through a packing fraction-controlled interplay between activity and particle interactions.
@@
This self-organization occurs through structural rearrangements mediated by topological defects, including the annihilation of positive and negative defects and the formation of large vortices. The number of defects, the mean squared displacement of the system, and the average curl of the displacement field together reveal coherent three-stage dynamics, highlighting the essential role of defects in the system’s evolution. From a kinematic perspective, defect annihilation and the creation of large-scale collective regions are associated with a turbulent-like inverse energy cascade, where energy transfers from small to large scales.
Our study demonstrates how large-scale collective motion emerges from a homogeneous ensemble of nonpolar Brownian particles, offering new insights into collective dynamics in biological and robotic systems, such as insect swarms, bird flocks, and robotic swarms.
@@ Acknowledgements.
Appendix A Methods
A.1 Identification of topological defects
@@
A.2 Characterization of structural rearrangements
@@
×$\times[r_{n}^{j}(t-\Delta t)-r_{0}^{j}(t-\Delta t)]\biggr{)}^{2},$
@@
×$\times[r_{n}^{j}(t-\Delta t)-r_{0}^{j}(t-\Delta t)],$
@@
The minimum value of $D(t,\Delta t)$ defines $D_{\text{min}}^{2}$ that quantifies the local deviation from affine deformation during a time interval $[t-\Delta t,t]$. $D_{\text{min}}^{2}$ is an excellent diagnostic for identifying local irreversible deformations and structural rearrangements.
Appendix B Spatial structure of topological defects
@@
Moreover, as shown in Fig. 7(a), the short range correlation between $+$ and -- TDs becomes more pronounced at late time, with the first peak in $g_{\pm}$ being much larger for long $\tau$. This suggests that, while the system relaxes towards the final non-equilibrium steady state, the TDs self-organize in short-range pairs and display a more ordered spatial distribution.
Appendix C Relaxation dynamics and dynamical heterogeinities
@@
where $q$ is the wave vector, $N$ is the total number of particles in the system, $r_{j}(t)$ is the position of particle $j$ at time $t$. Results in Fig.8(a) show that system exhibits progressively slower relaxation by increasing packing fraction.
[…]

### v2

@@ Introduction
To explore this physical mechanism further, we calculate $D_{\text{min}}^{2}$, which quantifies the strength of nonaffine particle displacements between two consecutive configurations (see Methods Methods for details) [80]. A large value for this quantity represents soft areas in the sense that structural relaxation is more intense than elsewhere. In Fig.4(a), we calculate the intensity of $D_{\text{min}}^{2}$ projected on each particle for $\tau=500s$ using a gray color map, with darker regions representing larger values of $D_{\text{min}}^{2}$. By visually comparing this color map with that presented in the third panel of Fig.3(a), it is clear that regions with large $D_{\text{min}}^{2}$ correspond to smaller values of $P_{c}$. This implies that areas with strong structural relaxation are those where motion is still random and poorly collective.
In Fig. 4(a), we represent the topological defects (TDs) with charges +1+1 and 1-1 using blue and red symbols, respectively. These TDs appear in pairs within regions of higher $D_{\text{min}}^{2}$ values. To quantify this observation, we define ‘soft spots’ (SS) as the top 5% of particles with the highest $D_{\text{min}}^{2}$ values and analyze their positional correlation with both positive and negative defects. The results are shown in Fig. 4(b) and (c). Each curve exhibits a peak at small $r$, with the peak height increasing slightly over time. This suggests a strong short-range correlation between soft spots and TDs. The correlation is slightly stronger for negative defects, though the difference is not substantial. This could be explained by the fact that TDs typically appear in pairs, which may balance out the correlation differences. Overall, the observations in Fig. 4 suggest that topological defects are involved in the system’s structural relaxation and contribute to the emergence of collective motion during out-of-equilibrium dynamics.
@@
We investigate the non-equilibrium dynamics of the system at $\phi=0.822$ across a full range of observation times, spanning six decades from $\tau=10^{-2}\,s$ to $\tau=10^{4}\,s$, and reveal a three-stage evolution that is evident in both topological characteristics, particle level dynamics, and collective motion. Three snapshots of the system dynamics in each of these stages have already been presented in Fig. 3 (see vertical red arrows in Fig. 5). An alternative representation highlighting the crossover between the first and second regimes is provided in the Supplementary Information.
@@
Figure 5(b) displays the mean square displacement (MSD), where the unit length is given in terms of the small particle diameter ( $D_{s}$). This quantity characterizes the system’s dynamics at the particle level. The MSD evolution follows three distinct stages. For $t<t_{1}$, the MSD exhibits sub-diffusive behavior, $\text{MSD}\propto t^{0.32}$, indicating that particle motion is constrained within a ‘cage’ on short timescales. Between $t_{1}$ and $t_{2}$, as the total number of TDs, $N_{\text{defects}}$, decreases more rapidly, particle dynamics accelerate while maintaining sub-diffusive characteristics with a different exponent, $\text{MSD}\sim t^{0.55}$. This regime coincides with the increasing mean cluster size of collective motion, as observed in the second snapshot of Fig. 3(a). Beyond $t_{2}$, particle dynamics slow as the system reaches a non-equilibrium steady state, where the MSD approaches a plateau. This indicates that at the highest packing fraction $\phi=0.822$, the system is in a dense liquid-like state, and the particle motion is very limited, i.e., frozen. This trend aligns with the stabilization of the total number of defects, consistent with defects potentially contributing to local relaxation. Ultimately, the system attains a collective steady state, emerging from the interplay between activity (energy injected through vibrations) and inelastic collisions.
@@
In conclusion, the number of defects, MSD, and average curl exhibit three distinct dynamical stages, establishing a link between topology, particle-level dynamics, and collective motion. These quantities are interrelated: at short time scales ($t<t_{1}$), particle motion is “trapped” due to frequent collisions, resulting in slow sub-diffusion in the MSD. Rapid fluctuations in the displacement field, driven by collisions, lead to efficient defect annihilation, causing a continuous decrease in defect numbers. Between $t_{1}$ and $t_{2}$, the system’s dynamics evolve more rapidly, driven by the formation of larger vortices in the displacement field. During this period, the defect number decreases at an accelerated rate. After $t_{2}$, the system reaches a steady state, where the number of defects, MSD, and average curl stabilize over time. The average curl attains its maximum value, signifying the emergence of stable large vortices. This accounts for the total topological number approaching a plateau and the onset of large-scale collective motion. In our system, defects are associated with local structural rearrangements that can contribute to the emergence of large vortices over long timescales. By analyzing the statistics of these topological defects, we obtain a simple and computationally inexpensive measure that captures the system’s three-stage dynamical behavior (see Fig. 5), showing excellent agreement with more traditional dynamical and coarse-grained observables.
Figure 6: Evolution of the average total topological charge with observation time $\mathbf{\tau}$. Different colors correspond to $\phi=0.773,0.811,0.822$. The vertical dashed lines indicates the characteristic timescales $t_{1},t_{2}$ for the highest packing fraction $\phi=0.822$.
To probe more deeply the topological nature of the active dynamics, we track the total topological charge, defined as the sum of the winding numbers of all defects in a single snapshot. In rigid systems with a well-defined order parameter and controlled boundaries, this charge is conserved in time. By contrast, in non-rigid fluids subject to activity and uncontrolled boundaries, the situation is more subtle.
We analyze three packing fractions, $\phi=0.822,0.811,0.773$, and compute the average total charge as a function of the observation timescale $\tau$. The results, averaged over $500$ realizations, are shown in Fig. 6.
At short times, below a critical scale $t^{*}\approx 10^{-1}$ s, the total charge remains approximately zero and conserved. This regime coincides with the solid-like behavior of our dense active fluid, previously linked to propagating shear waves and high-frequency rigidity [38].
At longer times, $t>t^{*}$, the system loses rigidity and charge conservation breaks down. For $\phi=0.822$ and $\phi=0.773$, the average charge stabilizes near unity at late times—remarkably close to the Euler characteristic of a 2D disk, which via the Poincaré–Hopf theorem fixes the total winding number to one. Since our system departs from the theorem’s assumptions (boundary conditions, activity, dissipation), whether this agreement is coincidental or physically meaningful remains to be understood.
By contrast, $\phi=0.811$ shows a steady growth of the average charge with time, pointing to a more intricate interplay between topology, activity, and dissipation, an avenue that calls for further exploration.
@@
In fluid turbulence, the inverse cascade is typically characterized in the inertial range through the third-order spatial correlation of velocity fluctuations, which reveals both the cascade process and its direction. For macroscopic active granular particles, however, the absence of a clear inertial range and the strongly dissipative nature of inelastic collisions and friction render this approach ineffective, as energy loss far exceeds any flux transferred through a cascade. We therefore adopt a different strategy: rather than attempting to quantify the exact flux, we focus on determining the direction of the cascade, which is experimentally more accessible.
Figure 7: Turbulent-like energy cascades as revealed by energy power spectra at $\phi=0.822$. Longitudinal (a) and transverse (b) kinetic energy spectra at different observation times $\tau$ as a function of the wave number $k$. The gray shaded background indicates the region dominated by the structure factor. (c) Power-law exponents $b$ for longitudinal ($b_{L}$) and transverse spectra ($b_{T}$). The background colors in panel (c) correspond to the system’s three dynamical stages shown in Fig.5.
@@
Nevertheless, a clear trend emerges in the intermediate and low range of $k$. In particular, at short times, both $E_{L}$ and $E_{T}$ grow with $k$, flattening in an intermediate $\tau$ range and then changing their slope at late times. To better characterize this behavior, we fit the various curves with a power-law ansatz $E_{L,T}\propto k^{-b_{L,T}}$. The extracted power-law exponent is presented in Fig. 7(c) as a function of the observation time $\tau$. Below $t_{1}$, we observe that $b_{L,T}$ are approximately identical and both negative, reflecting a spectrum dominated by small spatial scales, consistent with energy injection at the particle level by external vibrations. Around $t_{1}$, the spectrum starts to flatten, and the power-law exponents become slightly positive, presenting differences between the longitudinal and transverse sectors. Here $b_{L,T}=0$ denotes a flat spectrum with equal energy across different length scales, while $b_{L,T}>0$ signals energy transfer toward larger spatial scales. The evolution of $b_{L,T}$ with $\tau$ thus provides a clear indication of the cascade direction. Finally, above $t_{2}$, the power-law exponents reach a plateau value in approximately $b_{L}\approx 0.3$ and $b_{T}\approx 0.4$.
While Henkes et al. predicted $E(q)\sim q^{-1}\quad(b=1)$ [78], our measurements yield distinct exponents. This discrepancy can be attributed to the fact that, even at the highest packing fraction studied ($\phi=0.822$), the system remains in a liquid state rather than a jammed solid [38].
This kinematic behavior aligns well with the three-stage dynamics presented in Fig.5 and can be rationalized as follows. At short time scales, the energy injected by the active driving is gradually transferred to the individual particles, inducing a direct energy cascade towards microscopic scales. Around $t_{1}\approx 3\,s$, energy is evenly distributed across all scales, as indicated by the scaling exponent $b\approx 0$ shown in Fig.8(b). When the observation time exceeds 3s3\,s, energy starts transferring from small to larger scales, exhibiting a turbulent-like inverse cascade. This process becomes steady around $t_{2}\approx 60\,s$ when the energy transfer becomes roughly independent of time.
@@
To provide a self-consistent check of our independent measurement, we assumed that the mean squared vorticity (also known as enstrophy), $\overline{\Omega^{2}}$, follows $\overline{\Omega^{2}}\propto\left(\frac{\Omega_{0}}{\tau^{\alpha}}\right)^{2}$. This assumption draws an analogy with the turbulence cascade in 2D, where $\overline{\Omega^{2}}$ is conserved. Moreover, since the mean curl of the displacement field scales as $\Omega_{0}\propto\tau^{\alpha}$, this assumption removes the explicit time dependence from $\overline{\Omega^{2}}$.
We then performed a self-consistent check to determine whether this assumption aligns with the scaling of the displacement field, which itself depends nontrivially on $\tau$, namely ${\bf u}^{2}\propto\tau^{\beta}$. This time dependence can be estimated independently from the single-particle MSD curves in Fig. 5(b). Finally, we find that the results are consistent, as $\beta-2\alpha\approx 0$ for both $\tau<t_{1}$ and $t_{1}<\tau<t_{2}$.
From a physical perspective, this supports our initial assumption that the mean squared vorticity (enstrophy) is conserved, as expected for traditional two-dimensional turbulent flows.
@@
Our initial analysis focused on the highest packing fraction ($\phi=0.822$), where strong inelastic collisions balance the effects of activity, resulting in a late-time non-equilibrium steady state. Here, we extend the analysis to the lowest packing fraction ($\phi=0.309$), where the dynamics remain decorrelated and gas-like due to the weak inelastic interactions [38]. For low packing fraction, $\phi=0.309$, the system contains many empty domains. In this regime, the topological description is not well defined since the displacement vector cannot be reliably approximated by a continuous field. Nevertheless, for completeness, we also present the results for $\phi=0.309$ in Fig. 8(a).
Fig.8(a) shows the evolution of the total number of topological defects, $N_{\text{defects}}$, for four packing fractions ($\phi=0.309$, $0.618$, $0.773$, and $0.811$). At a high packing fraction ($\phi=0.811$), the behavior of $N_{\text{defects}}$ resembles that observed at the maximal packing fraction shown in Fig. 5(a).
@@
In Fig.8, the strongest nonmonotonic behavior appears at $\phi=0.618$ and $\phi=0.773$, both in the defect number and in the scaling exponents $b_{T,L}$. An energy spectrum with $b_{T,L}\approx 0$, corresponding to nearly equal energy across length scales, is highly unusual and has previously been linked to uncorrelated bacterial swimmers in the dilute limit[93, 94]. In our system, however, this behavior originates from transient domains of collective motion that coexist with disordered regions and rapidly reorganize, as illustrated by the displacement field at $\phi=0.773$, $\tau=1\,\mathrm{s}$ (see Movie S3.mp4 in the SI Movies). Such spatiotemporal mixtures generate flat spectra and give rise to the nonmonotonic trends in Fig. 6. This interpretation is consistent with the sharp increase of the structural relaxation time around $\phi\approx 0.773$ (see Fig.S5 in the Supplementary Information), which indicates glassy dynamics.
@@ Conclusions
Using active granular vibrators as a model for apolar dry active Ornstein–Uhlenbeck particles, we conducted a comprehensive experimental study to investigate the topological features that give rise to steady-state large-scale collective dynamics and turbulent-like inverse energy cascades. The dynamical behavior of the system evolves through a packing fraction-controlled interplay between activity and particle interactions.
@@
This self-organization occurs through structural rearrangements associated with topological defects, including the annihilation of positive and negative defects and the formation of large vortices. The number of defects, the mean squared displacement of the system, and the average curl of the displacement field together reveal coherent three-stage dynamics, highlighting the notable role of defects in the system’s evolution, which can serve as an efficient and practical way to quantify the system’s behavior. From a kinematic perspective, defect annihilation and the creation of large-scale collective regions are associated with a turbulent-like inverse energy cascade, where energy transfers from small to large scales.
Our study demonstrates how large-scale collective motion emerges from a homogeneous ensemble of apolar dry active Ornstein–Uhlenbeck particles, offering new insights into collective dynamics in biological and robotic systems, such as insect swarms, bird flocks, and robotic swarms.
@@ Acknowledgements.
Methods
.1 Single-particle correlation function
For single-particle dynamics (Fig. 2), the velocity autocorrelation function is defined as
$C_{v}(t)=\left\langle\frac{\mathbf{V}(0)\cdot\mathbf{V}(t)}{\mathbf{V}(0)\cdot\mathbf{V}(0)}\right\rangle,$ (10)
where $\mathbf{V}(t)=\frac{\mathbf{r}(t+\delta t)-\mathbf{r}(t)}{\delta t}$ with $\delta t=0.025\,\mathrm{s}$. The orientation correlation function is given by
$C_{n}(t)=\left\langle\cos[\theta(t)-\theta(0)]\right\rangle.$ (11)
Since the particles are non-polar, $\theta$ is defined by a short marker line on each particle to enable tracking of rotational motion (Fig. 1(a)).
.2 Identification of topological defects
@@ Methods
.3 Characterization of structural rearrangements
@@
$\times[r_{n}^{j}(t-\Delta t)-r_{0}^{j}(t-\Delta t)]\biggr{)}^{2},$
@@
$\times[r_{n}^{j}(t-\Delta t)-r_{0}^{j}(t-\Delta t)],$
@@
The minimum value of $D(t,\Delta t)$ defines $D_{\text{min}}^{2}$ that quantifies the local deviation from affine deformation during a time interval $t-\Delta t,t]$. $D_{\text{min}}^{2}$ is an excellent diagnostic for identifying local irreversible deformations and structural rearrangements.
Supplementary Information
Measurement of restitution coefficient
To measure the restitution coefficient, particle A was fixed sideways on the substrate, while particle B, also placed sideways, was released from a height $h_{1}=20\,\mathrm{cm}$ directly above A. After the head-on side-surface collision, the rebound height $h_{2}$ of particle B was recorded with a high-speed camera. Repeated trials with different particle pairs yielded the average rebound height, from which the restitution coefficient was obtained as
$\epsilon=\frac{|v_{2}|}{|v_{1}|}=\frac{\sqrt{2gh_{2}}}{\sqrt{2gh_{1}}}=\sqrt{\frac{h_{2}}{h_{1}}},$ (16)
where $v_{1}$ and $v_{2}$ are the pre- and post-collision velocities, and $g$ is the gravitational acceleration.
Single particle dynamics
In our experiments, the persistence time $\tau_{p}$ is defined as the cutoff time marking the end of the initial superdiffusive regime. Specifically, $\tau_{p}$ corresponds to the time point at which the log-log mean squared displacements (MSD) slope first decreases below 1 after exceeding it:
$k^{*}=\frac{d(\log(\mathrm{MSD}))}{d(\log(t))},\qquad\tau_{p}=\max\left(t(k^{*}>1)\right).$ (17)
It is worth noting that $\tau_{p}$ is closely related to the packing fraction $\phi$ due to interparticle interactions, as the frequency of interparticle collisions governs the persistence time. At the packing fraction of $\phi=0.822$, the MSD of large and small particles are presented in the Fig. 9.
From this analysis, we can obtain the local slope of the MSD for large and small particles respectively in the log-log axis $k^{*}$, as shown in Fig. 10.
At this high packing fraction, the particles do not exhibit superdiffusive behavior within the temporal resolution of our experiment ($\Delta t=0.025\,\mathrm{s}$). We therefore approximate the persistence time of both large and small particles to correspond with the temporal resolution limit of $0.025\,\mathrm{s}$ under these conditions. Thus, we conclude that
$\tau_{p}=0.025\,\mathrm{s},$ (18)
for both large and small particles.
In the following, we introduce the calculation of the persistence length and the Péclet number. The particles’ diameter is $D_{l}=22.4\,\mathrm{mm}$ for large particles and $D_{s}=16\,\mathrm{mm}$ for small particles. For both type of particles we estimate the root mean square active velocity $v_{0}$ from single-particle experiments:
$v_{0,l}=\sqrt{\langle\overline{\mathbf{v}^{2}(t)}\rangle}\approx 1.58\,D_{s}\cdot\mathrm{s}^{-1}\approx 25.3\,\mathrm{mm}\cdot\mathrm{s}^{-1},$ (19)
$v_{0,s}=\sqrt{\langle\overline{\mathbf{v}^{2}(t)}\rangle}\approx 1.43\,D_{s}\cdot\mathrm{s}^{-1}\approx 22.9\,\mathrm{mm}\cdot\mathrm{s}^{-1}\,\,.$ (20)
Here, the subscripts $l$ and $s$ respectively represent large and small particle. The notation $\langle\cdots\rangle$ denotes an average over repeated experimental realizations, while the overline indicates a time average along individual particle trajectories.
Then, the persistence length $l_{p}$ is
$l_{p,l}=v_{0,l}\cdot\tau_{p}=25.3\,\mathrm{mm/s}\times 0.025\,\mathrm{s}=0.6325\,\mathrm{mm},$ (21)
$l_{p,s}=v_{0,s}\cdot\tau_{p}=22.9\,\mathrm{mm/s}\times 0.025\,\mathrm{s}=0.5725\,\mathrm{mm}\,.$ (22)
Finally, the Péclet number $\mathrm{Pe}$ is
$\mathrm{Pe}_{l}=\frac{l_{p,l}}{D_{l}}=25.3\,\mathrm{mm/s}\times 0.025\,\mathrm{s}\times\frac{1}{22.4\,\mathrm{mm}}=0.0282,$ (23)
$\mathrm{Pe}_{s}=\frac{l_{p,\mathrm{s}}}{D_{\mathrm{s}}}=22.9\,\mathrm{mm/s}\times 0.025\,\mathrm{s}\times\frac{1}{16\,\mathrm{mm}}=0.0358,$ (24)
where the subscripts $l$ and $s$ represent large and small particle respectively.
Discussion of the effect of system boundaries on the final vortex
The flower-shaped boundary is expected to have a negligible impact on the observed vortex field. This is supported by the observation that, when large internal vortices form, the displacements of boundary particles remain small and exhibit disordered directions, indicating minimal coupling with the internal vortex motion.
For comparison, in a monodisperse system at high packing fraction, where crystallization occurs, the boundary geometry can influence flocking motion due to long-range translational order. In contrast, in our bidisperse, structurally disordered system, such boundary effects are expected to be negligible.
Spatial structure of topological defects
@@
Moreover, as shown in Fig. 11(a), the short range correlation between ++ and - TDs becomes more pronounced at late time, with the first peak in $g_{\pm}$ being much larger for long $\tau$. This suggests that, while the system relaxes towards the final non-equilibrium steady state, the TDs self-organize in short-range pairs and display a more ordered spatial distribution.
Relaxation dynamics and dynamical heterogeinities
@@
where $q$ is the wave vector, $N$ is the total number of particles in the system, $r_{j}(t)$ is the position of particle $j$ at time $t$. Results in Fig. 12(a) show that system exhibits progressively slower relaxation by increasing packing fraction. The structural relaxation time $\tau_{a}$ as a function of packing fraction $\phi$ is presented in Fig. 13. Here, $\tau_{a}$ is defined via the self-intermediate scattering function as the time at which it decays to $1/e$.
The four-point susceptibility in Fig. 12(b) describes the degree of dynamical heterogeneity in space within a time interval, and is defined as: $\chi_{4}(t)=N[\langle q_{s}(t)^{2}\rangle-\langle q_{s}(t)\rangle^{2}]$ , where $q_{s}(t)=(1/N)\sum_{i=1}^{N}w(|r_{\mathrm{i}}(t)-r_{\mathrm{i}}(0)|)$ , $w=1(0)$ if $|r_{\mathrm{i}}(t)-r_{\mathrm{i}}(0)|<(>)0.5D$. $N$ is the total number of particles in the system, $r_{i}(t)$ is the position of particle $i$ at time $t$, and $D$ is the average particle diameter. The average $\langle\cdot\rangle$ is performed over all particles and all initial times. The peak position in $\chi_{4}(t)$ indicates the time during which the dynamics exhibit the greatest heterogeneity, while the peak height reflects the intensity of these heterogeneities. As the packing fraction increases, the time at which the system reaches its maximum dynamical heterogeneity is delayed. Meanwhile, from a packing fraction of $\phi=0.309$ to $0.811$, the peak intensity of dynamical heterogeneity also increases. Due to experimental limitations, the data for $\phi=0.822$ do not exhibit a distinct peak within the accessible observation time. However, we speculate that a peak would emerge over a longer time scale.
Characterization of three stage non-equilibrium dynamics
To highlight the transition at $t=t_{1}$, we fitted the data for $t<t_{1}$ with a power-law function. As shown in Fig. 14, while the data follow the fit closely for $t<t_{1}$, clear deviations arise for $t>t_{1}$, indicating a distinct change in the dynamical behavior.
Figure 9: Mean square displacements (MSD) for large and small particles at packing fraction $\phi=0.822$, in units of the square of the small particle diameter $D_{s}^{2}$. The background colors correspond to the three dynamical stages shown in Fig. 4 of the main text.
Figure 10: The local slope $k^{*}$ of the mean square displacements (MSDs) curves in Fig. 9.
@@
Figure 13: Structural relaxation time versus packing fraction $\phi$. The structural relaxation time is determined from the self-intermediate scattering function as the time at which it decays to $1/e$.
Figure 14: Characterization of the three-stage non-equilibrium dynamics. (a) Topological signatures. (b) Particle-level dynamics. (c) Collective motion. In all panels the solid blue line is a power-law fit of the experimental data in the range $t<t_{1}$ extended up to $t=t_{2}$. The deviation of the experimental data from this power-law behavior at $t_{1}$ is evident in all three physical quantities.